# Detection of Synthetic Face Images: Accuracy, Robustness, Generalization

### v1

@@ Abstract
An experimental study on detecting synthetic face images is presented. We collected a dataset, called FF5, of five fake face image generators, including recent diffusion models. We find that a simple model trained on a specific image generator can achieve near-perfect accuracy in separating synthetic and real images. The model handles common image distortions (reduced resolution, compression) by using data augmentation. Moreover, partial manipulations, where synthetic images are blended into real ones by inpainting, are identified and the area of the manipulation is localized by a simple model of YOLO architecture. However, the model turned out to be vulnerable to adversarial attacks and does not generalize to unseen generators. Failure to generalize to detect images produced by a newer generator also occurs for recent state-of-the-art methods, which we tested on Realistic Vision, a fine-tuned version of StabilityAI’s Stable Diffusion image generator.
@@ 1 Introduction
Image synthesis has made remarkable progress in recent years, thanks to the advances of generative models such as Generative Adversarial Networks (GANs) [22] and Diffusion Models [34]. Synthesized images are becoming increasingly realistic and hardly distinguishable from real ones to the naked eye of an average human and even of an expert, see Fig. 1. However, this progress also poses serious threats to individuals and society, as synthesized images, a.k.a. ‘deep fakes’ [29], can be used and have already been used for malicious purposes, such as fake porn [26], fake video calls [3, 39], fake news [42], or fake videos in election campaigns [32, 28]. Therefore, it is important to develop effective and robust methods to detect and expose fake images, especially in the domain of faces, which are often the target of the attacks.
In this paper, we present an experimental study in which we show some key properties of neural fake-face detectors. We use a standard classifier, and we do not primarily aim to push accuracy on standard datasets, but rather focus on other aspects of the problem. In particular, we investigate the generalization ability of the detector to unseen generators, the robustness to various image degradations and input sizes, the vulnerability to adversarial attacks, and the localization of artifacts in the case of partial manipulation of real images.
@@
In addition to detectors which are accurate in spotting recent generator images (namely, the Stable diffusion [34] –- Realistic Vision [7]), our main contribution is a thorough analysis of forgery detectors that revealed many intriguing properties. Our contributions are summarized below.
@@
    Cross-generator-detector testing. We show that while it is surprisingly easy to train a detector for a specific synthetic image generator, its accuracy drops dramatically when tested on images produced by a different generator, which was not trained for. This effect is partially reduced by training the detector on images generated by multiple different generators. Then, the detector generalizes better to unseen generated images. We quantify this effect and show learning curves that demonstrate the accuracy as a function of a number of training images.
@@
    Robustness to input size and degradations (blur, compression) analysis. We tested the detector against blur (reflecting input resolution), JPEG compression, and by masking the input image. The detector is surprisingly robust to the degradations and is even more robust when the degradations are added to the training set as data augmentation. For instance, our detector is able to spot synthetic images from a patch as small as 25$\times$25 px with accuracy about 70%.
@@
    Adversarial attacks vulnerability investigation. We demonstrate that adversarial images can be easily found and fool the detector to classify the synthetic image as real. Moreover, we show that residuals found for a particular fake image have the adversarial effect on other images and can also fool a different model of a very different architecture. We tested convolutional networks and a vision transformer.
@@
The rest of the paper is organized as follows. The related work is summarized in Sec. 2, the experimental analysis is presented in Sec. 3, and finally, Sec. 4 gives conclusions.
@@ 2 Related Work
Forensic low-level signal detectors are another class of methods. They exploit the spectral signatures of synthetic images. Inspiration probably came from the forensic recognition of a camera device [5]. More recently, researchers discovered that the residual spectra of synthetic images contain typical anomalies, which creates a spectral fingerprint of a synthetic generator [43, 8, 9]. A simple method based on the spectral fingerprint [12] reports a high detection accuracy. However, these methods may not be robust to image transformations, such as resizing or compression, that can alter the signal characteristics.
In a similar spirit, other more recent methods suggested that the information for fake image detection is deeply embedded in the image signals and can be detected independently of the image content. Work [4] proposed a method that uses a convolutional neural network with a narrow receptive field to detect fake images from signal patches, and showed that some regions, such as hair, are more discriminative than others.
Another paper [43] has shown that the detection of synthetic images generated by GANs is relatively easy, as they exhibit certain distinctive features and a standard CNN classifier can be easily trained to capture this difference. The paper even shows a good generalization ability, in case the model is tested on images produced by a generator for which it was not trained for. However, they used only GAN-based generators. In this paper we will show that the generalization of detectors to unseen generators between recent models (especially of different architectures GANs/diffusion models) is poor. We show that recent state-of-the-art synthetic image detectors fail completely when tested on images produced by novel unseen generators.
Besides detection, some recent works have also addressed the localization of fake images, which aims to segment the manipulated areas of the real images. The problem is challenging considering possibly a small area of manipulation. Some methods do not use any special architecture for localization, but rely on post-processing techniques. Recent paper [38] compares a popular Grad-CAM [35] to highlight the regions that contribute to the classification decision and the scanning technique of [4] to localize synthetic regions in partially manipulated images. Other methods use more complex architectures, such as multi-branch network [17], or dense self-attention network [18], to explicitly learn the localization maps. Paper [25] fine-tunes a large segmentation model (SAM) [24] to adapt it to the fake image domain. We show that precise localization results are achieved for relatively small regions using a simple YOLO-based architecture [31], namely YOLOv8 [41], as long as the fake images are composed of images produced by the same generator model that we trained on.
@@ 3 Experimental Analysis
For the negative class of real images, we use images from the FFHQ dataset [21]. All synthetic and real images underwent the same preprocessing procedure, cropping with the same margin, aligning using facial landmarks, and resampling to $224\times 224$ px.
@@ 3.1 Cross-generator testing
Table 1: Cross-generator testing. Each cell $(row,col)$ shows test accuracy in percent of the models trained (a) on generator $row$ / (b) all without generator $row$ , tested on generator $col$.
@@
(a) Training on a single generator (b) Leave one out training
@@
We can see in Tab. 1a that if the detector is trained on the same model as it is tested (diagonal of the table), the accuracy is perfect for RV5.1, SD 1.4, PG, and very high for SG. The accuracy is only 80% for FA. FA, FaceApp [14], a commercial app with unknown technology behind, probably blends the real face with some manipulations, making it harder to identify. However, we can clearly see (off the diagonal) that accuracy drops close to chance when we test on images produced by models for which the detector was not trained on. Interestingly, this is not the case of FA, which achieves even higher accuracy on PG, which might indicate similar technology, but the converse is not true. The generalization does not occur for even very similar models, the diffusion models RV5.1 and SD 1.4 share the same architecture. The first is a fine-tuned version of the latter.
In Tab. 1b, when the detector is trained on multiple models, a certain level of generalization to unseen generators is achieved for some models, as seen in the diagonal now. SD1.4 seems to generalize well while it was not trained on it. However, RV5.1 is similar to SD1.4, but the generalization is not reciprocal. PG seems to generalize too, and the rest is close to chance.
@@
Interestingly, for some generators, only units or small tens of training samples are needed to significantly improve detection accuracy.
@@
Model Wang 2020 [43] HiFi 2023 [17] Durall 2019 [12] ResNET-50 (ours)
Accuracy 48.3 44.2 87.7 99.5
Comparison with the state of the art.
We evaluate recent fake image detectors on our test set produced by RV5.1 generator. We tested Wang 2020 [43], HiFi 2023 [17], which both provide pre-trained models, and Durrall 2019 [12] which we trained on the independent training split of the RV5.1 dataset by using the training script provided by the authors. Our ResNET-50 was trained on the same set.
The results are shown in Tab. 2. Although Wang 2020 [43] claimed to generalize to unseen generators, it no longer holds with novel generators – the diffusion model RV5.1. The model achieved accuracy close to chance. It was trained on GAN-based generators which does not provide any generalization to the recent diffusion RV1.5 generator. Moreover, the recent HiFi 2023 [17] also failed, despite being trained on diffusion models. It achieves low accuracy 44.2% as it confused almost all RV5.1 images with real ones, while classifying some of the real images as synthetic. This generalization failure is particularly striking, as this method presented in the last CVPR conference is supposed to be the state of the art. This finding reveals the difficulty of the problem of detecting unseen generators.
Durral 2019 [12] resulted in an inferior accuracy compared to our model. The reason is probably that it uses very simple features, magnitude spectrum radius and logistic regression.
@@ 3.2 Detector accuracy for input degradation
The results are shown in Fig. 3. We can see that the detector proves a good robustness to the degradations, especially for the second scenario with re-training. For instance, the detector achieves accuracy about 80% for Gaussian blur $\sigma=17$ px, 90% for JPEG quality 10, and 70% for a patch as small as 25$\times$25 px.
@@ 3.3 Adversarial attacks
In this section, we study the vulnerability of our detector to adversarial attacks. An adversarial attack means performing a hardly perceptible modification of an image (residue) that causes a change in classification of the detector, i.e., a synthetic image is classified as real. First, we find the adversarial residua for a given image. We use the simple fast gradient sign method (FGSM) [16]. The adversarial residuum is then an image of the same size as the input receptive field of the model that contains signs of the gradient of the output class score with respect to the input pixel intensities in each channel. This is easily calculated by backpropagation without any optimization.
We tested the attacks on three different architectures of detectors: ResNET-50 [19], Xception [6], ViT-tiny [40], which we trained on our training set RV5.1. Two of the models is of perfect 100% accuracy on the test set, ViT-tiny achieved 83%.
The resulting residua are shown in Figs. 5 and 6. The residua are scaled up multiple times in the figures, otherwise the pattern would not be visible. All these residua scaled by strength $\epsilon$ if summed with the original images will switch the classification of the corresponding detector to “real". Residua for different models appear different and the pattern is visibly influenced by the structure of the original image, as seen in Fig. 6.
We then tested whether the residua found for a specific image and a specific detector act adversarially on another image for the same or a different detector model. We measured the success of attacks by the confusion rate, which depicts a percentage of test cases when the model switched the classification due to the attack from “fake” to “real” over the number of “fake” decisions prior to the attack.
First, we performed the cross-sample test, where the adversarial residua are found for a given image and are tested also on different images. This was done for the three models separately. The confusion rates for increasing strength of the residua $\epsilon$ are presented in Tab. 3. Cross-sample attacks are seen to work and are more successful with greater strength $\epsilon$. We show this experiment as an interesting insight, although this scenario is not so relevant practically, since an attacker would prepare an optimum residue right for his fake image.
Much more important is the cross-model scenario, where the adversarial residua are found for a given image and a given model, and are tested also on other models of different architectures. The results for increasing strength of the residua $\epsilon$ are summarized in Tab. 4.
@@
Table 3: Cross-sample adversarial attacks. Adversarial residua found for a given image were tested on different images. Table show confusion rate for increasing strength of residua $\epsilon$. The experiment was done separately for three model architectures.
$\epsilon=0.01$ $\epsilon=0.02$ $\epsilon=0.05$
ResNET-50 2.20 67.40 100.00
Xception 96.34 100.0 100.00
ViT-tiny 8.63 25.88 97.25
Table 4: Cross-architecture adversarial attacks for increasing strength of the residua $\epsilon$. Each cell $(row,col)$ corresponds to confusion rate in percent. The attack was targeted against model of architecture in $row$ and tested against the model of architecture in $col$.
@@
Figure 5: Adversarial images crafted to confuse the classification model. They are created by summing the original (top left) image with adversarial residue (top row) scaled by $\epsilon$. Results shown for three models (ResNET-50, Xception, ViT-tiny) with increasing strength $\epsilon$.
Figure 6: Other examples of adversarial residua shown for input images and three fake detector models (ResNET-50, Xception, ViT-tiny).
@@ 3.5 Implementation details
We used AUTOMATIC1111’s Stable Diffusion Web UI [2] for our experiments. Besides graphical interface with many plugins, it provides a convenient batch processing.
To generate our dataset (RV5.1 and SD1.4), we used dynamic prompt [13]. This is the Web UI extension that implements an expressive template language for the generation of random or combinatorial prompts. In particular, we used the following prompt to get diversity in our datasat: “RAW photo, {older |||| younger} {man |||| woman |||| lady |||| girl |||| boy} { {smiling |||| staring} |||| with glasses |||| with hat |||| with {brown |||| blonde |||| dark} {straight |||| curly |||| short} hair }, high quality portrait taken with Nikon camera, in {nature |||| a city |||| a room |||| an office |||| a park |||| a street |||| a forest }”. All models were trained with PyTorch framework.

### v2

@@ Abstract
An experimental study on synthetic face image detection is presented. We introduce FF5, a dataset of five fake face generators, including recent diffusion models. A baseline model trained on a specific generator achieves near-perfect accuracy in distinguishing synthetic from real images and handles common distortions (e.g., compression) via data augmentation. Additionally, partial manipulations, where synthetic content is blended into real images, can be detected and localized using a YOLO-based model. However, the model is vulnerable to adversarial attacks and fails to generalize to unseen generators – a limitation shared by state-of-the-art methods. Testing on Realistic Vision, a fine-tuned version of Stable Diffusion, confirms these challenges. Our study provides a quantitative evaluation of key properties and empirical evidence that deepfake detectors primarily learn generator fingerprints embedded in the signal.
Keywords:
Deepfake face generated images detection localization.
@@ 1 Introduction
Image synthesis has made remarkable progress in recent years, thanks to the advances of generative models such as Generative Adversarial Networks (GANs) [24] and Diffusion Models [41]. Synthesized images are becoming increasingly realistic and hardly distinguishable from real ones to the naked eye of an average human and even of an expert, see Fig. 1. However, this progress also poses serious threats to individuals and society [18, 25], as synthesized images, also known as ‘deep fakes’ [33], can be used for malicious purposes, such as fake porn [29], fake video calls [2, 48], fake news [52], or fake videos in election campaigns [39, 31]. Therefore, it is important to develop effective and robust methods to detect and expose fake images, especially in the domain of faces, which are often the target of the attacks.
In this paper, we present a comprehensive experimental study that uncovers key properties of neural fake-face detectors. Rather than solely optimizing accuracy on standard datasets, we take a broader approach, using models with standard architectures to explore fundamental challenges in synthetic image detection. Specifically, we investigate the generalization ability of detectors when faced with unseen generators, their robustness to various image degradations and input sizes, their vulnerability to adversarial attacks, and their effectiveness in localizing manipulated regions within partially altered real images.
@@
In addition to detectors which are accurate in spotting recent generator images (namely, the Stable diffusion [41] –- Realistic Vision [6]), our main contribution is a thorough analysis of forgery detectors that revealed many intriguing properties. To the best of our knowledge, no existing work offers a similar in-depth analysis within a compact and easily reproducible setup. Our contributions are summarized below.
@@
    Cross-generator-detector testing. We show that while it is surprisingly easy to train a detector for a specific synthetic image generator, its accuracy drops dramatically when tested on images produced by a different generator, which was not trained for. This effect is not much reduced by training the detector on images generated by multiple different generators. We quantify this effect and show learning curves that demonstrate the accuracy as a function of a number of training images.
@@
    Robustness to input size and degradations analysis. We tested the detector against blur (reflecting input resolution), JPEG compression, and input patch size reduction via masking. It demonstrated strong robustness, further improving when degradations were included as data augmentation. Notably, our detector spots synthetic images from a $\times25$ px patch with about 70% accuracy.
@@
    Adversarial attacks vulnerability investigation. We demonstrate that adversarial images can easily be found to deceive the detector to classify synthetic image as real. Moreover, we show that residuals found for a particular detector model can also fool other models of a very different architecture. We tested both convolutional networks and a vision transformer.
@@
The rest of the paper is organized as follows. The related work is summarized in Sec. 2, the proposed methodology and results of the experimental analysis are presented in Sec. 3, and finally, Sec. 4 gives conclusions.
@@ 2 Related Work
Forensic low-level signal detectors are another class of methods. They exploit the spectral signatures of synthetic images. Inspiration probably came from the recognition of a camera device [4]. More recently, researchers discovered that the residual spectra of synthetic images contain typical anomalies, which creates a spectral fingerprint of a synthetic generator [53, 7, 8]. A frequency domain method is presented in [12].
In a similar spirit, other methods suggested that the information for fake image detection is deeply embedded in the image signals and can be detected independently of the image content. Chai et al. [3] used a CNN with a narrow receptive field to detect fake images from signal patches, highlighting hair as a key discriminative region. Tan et al. [46] proposed to spot upsampling artifacts by modeling neighboring pixel relationships. Shiohara et al. [43] proposed detecting blending artifacts in deepfake images.
Wang et al. [53] showed that GAN-generated images have distinctive features, making detection easy with a CNN classifier. Their model generalized well to unseen generators but was tested only on GAN-based images. Very promising approach to detect AI-generated images is by using CLIP [36], as a powerful image encoder, followed by a lightweight classifier head [34, 9, 56]. The authors report promising generalization abilities.
In this paper, we demonstrate that, in the leave-one-out setup, the generalization of detectors to unseen generators is poor. We show that even recent state-of-the-art synthetic image detectors either fail completely or perform low when tested on images produced by novel unseen generators. A recent work related to ours [35] tested generalization by learning on samples from large scale dataset comprising 4.8k distinct synthetic generator models.
Besides detection, some recent works have also addressed the localization of fake images, which aims to segment the manipulated areas of the real images. The problem is challenging considering possibly a small area of manipulation. Some methods do not use any special architecture for localization, but rely on post-processing techniques. Recent paper [47] compares a popular Grad-CAM [42] to highlight the regions that contribute to the classification decision and the scanning technique of [3] to localize synthetic regions in partially manipulated images. Other methods use more complex architectures, such as multi-branch network [17], or dense self-attention network [19], to explicitly learn the localization maps. Paper [28] fine-tunes a large segmentation model (SAM) [27] to adapt it to the fake image domain. We show that precise localization results are achieved for relatively small regions using a simple YOLO-based architecture [38], namely YOLOv8 [51], as long as the fake images are composed of images produced by the same generator model that was trained on.
@@ 3 Methodology and Experimental Results
For the negative class of real images, we use images from the FFHQ dataset [23]. All synthetic and real images underwent the same preprocessing procedure, aligning using facial landmarks, cropping with the same margin, and resampling to $224\times 224$ px.
@@ 3.1 Cross-generator testing
Table 1: Cross-generator testing. Each cell $(row,col)$ shows test accuracy in percent of the models trained (a) on generator rowrow / (b) all without generator rowrow, tested on generator colcol.
@@
(a) Training on a single generator (b) Leave-one-out training
@@
We can see in Tab. 1a that if the detector is trained on the same model as it is tested (diagonal of the table), the accuracy is perfect for RV5.1, SD 1.4, PG, and very high for SG. The accuracy is only 80% for FA. FA, FaceApp [14], a commercial app with unknown technology behind, probably blends the real face with some manipulations, making it harder to identify. However, we can clearly see (off the diagonal) that accuracy drops close to chance when we test on images produced by models for which the detector was not trained on. Interestingly, this is not the case of FA, which achieves even higher accuracy on PG, which might indicate similar technology, but the converse is not true. The generalization does not occur for even very similar models, the diffusion models RV5.1 and SD 1.4 share the same architecture.
In Tab. 1b, when the detector is trained on multiple models, a certain level of generalization to unseen generators is achieved for some models, as seen in the diagonal now. SD1.4 seems to generalize well while it was not trained on it. However, RV5.1 is fine-tuned version of SD1.4, but the generalization is not reciprocal. PG seems to generalize partially as it is another GAN as SG. The rest is close to chance.
Note that the cross-dataset experiment includes several generators representing the fake class, while the real class is represented by a single source, the FFHQ dataset. This is a limitation, as real-world face images exhibit significantly more diversity than what is captured by this dataset. Therefore, these results should be interpreted as an optimistic upper bound; accuracy is expected to decline when a more diverse and previously unseen real dataset is used.
@@
Interestingly, the learning curves are steep. For some generators, only a few units or small tens of training samples are sufficient to significantly improve detection accuracy, indicating that the model quickly captures the fingerprint of the new generator.
@@
Comparison with the state of the art.
We evaluate recent fake image detectors on our test set produced by the RV5.1 generator, see Tab. 2. The first four methods [53, 17, 46, 9] provide pre-trained models, while the last two models, Durall [12] and our ResNET-50, were trained on an independent training split of the RV5.1 dataset.
Wang [53] claims to generalize to unseen generators, but this does not hold for novel diffusion models such as RV5.1. The model achieved accuracy close to chance, likely because it was trained on GAN-based generators, which do not generalize to the recent diffusion-based RV5.1 generator. HiFi [17] failed despite being trained on diffusion models. Tan [46] performed slightly above chance level, even though the paper reports a generalization to unseen generators by spotting upsampling artifacts. Cozzolino [9] achieves better, but still low, accuracy despite being trained on Stable Diffusion and reporting generalization abilities via CLIP [36].
@@
Model Wang [53] HiFi [17] Tan [46] Cozzolino [9] Durall [12] ResNET-50
CVPR’20 CVPR’23 CVPR’24 CVPRw’24 CVPR’20 ours
Accuracy 48.3 44.2 64.1 70.0 87.7 99.5
Our simple ResNET-50, when trained on the RV5.1 training split, achieved near-perfect recognition. In contrast, Durall [12] resulted in inferior accuracy, likely due to its reliance on very simple features – magnitude spectrum radius and logistic regression.
This experiment demonstrates that generalization to unseen generators remains an unsolved problem in practice. A trivial classifier, when trained on examples from the target generator, outperforms more sophisticated methods. The likely reason is that detectors overfit to known generator fingerprint and are unable to identify more universal traces that separate synthetic and real samples.
The remaining experiments are conducted using our models trained within the dataset, since the competing methods do not generalize well and some perform at chance level, as seen in Tab. 2.
@@ 3.2 Detector accuracy for input degradation
The results are shown in Fig. 3. We can see that the detector proves a good robustness to the degradations, especially for the second scenario with re-training. For instance, the detector achieves accuracy about 80% for Gaussian blur $\sigma=17$ px, 90% for JPEG quality 10, and 70% for a patch as small as $\times25$ px.
These findings corroborate that the fingerprint is strong, survives severe image degradations, and can be identified through discriminative learning. The small-patch experiment proves that the fingerprint exists at a low (signal) rather than a high (semantic) level.
@@ 3.3 Adversarial attacks
In this section, we study the vulnerability of our detector to adversarial attacks. An adversarial attack means performing a hardly perceptible modification of an image (residue) that causes a change in classification of the detector, i.e., a synthetic image is classified as real. First, we find the adversarial residua for a given image. We use the fast gradient sign method (FGSM) [16]. The adversarial residuum is then an image of the same size as the input receptive field of the model that contains signs of the gradient of the output class score with respect to the input pixel intensities in each channel. This is easily calculated by backpropagation without any optimization.
We tested the attacks on three different architectures of detectors: ResNET-50 [20], Xception [5], ViT-tiny [49], which we trained on our training set RV5.1. All models achieved a perfect 100% accuracy on the test set.
The resulting residua are shown in Fig. 5. The residua are scaled up multiple times in the figures, otherwise the pattern would not be visible. All these residua scaled by strength $\epsilon$ if summed with the original images will switch the classification of the corresponding detector to “real". Residua for different models appear different and the pattern is visibly influenced by the structure of the original image, as seen in Fig. 5.
We measured the success of attacks by the confusion rate, which depicts a percentage of test cases when the model switched the classification due to the attack from “fake” to “real” over the number of “fake” decisions prior to the attack.
We tested the cross-model scenario, where the adversarial residua are found for a given image and a given model, and are tested also on other models of different architectures. The results for increasing strengths of the residua, $\epsilon$, are summarized in the tables in Fig. 5.
@@
A defense against adversarial attacks on deepfake detectors should be implemented in practice. It has been reported, e.g., in [32], that image compression or low-pass filtering can mitigate adversarial effects. However, as we have shown in this paper, such techniques also weaken the deepfake generator fingerprint. Another option is to use adversarial training [57] or detectors for adversarial patterns, though this remains a non-trivial problem [50], similar in spirit to deepfake detection. Therefore, we believe these problems should be studied together in future research.
@@
Figure 5: Cross-architecture adversarial attacks for increasing strength of the residua $\epsilon$. Left: Examples of adversarial residua for specific models created with FGSM method. Right: Results showing confusion rate in percent for each cell $(row,col)$. The attack was targeted against model of architecture in rowrow and tested against the model of architecture in colcol.
@@ 3.4 Localizing partial manipulations
On the other hand, we attribute the success of the detector trained for this particular manipulation technique to its sensitivity in identifying small patches of the synthetic signal, as discussed in Sec. 3.2, and possibly to its ability to detect subtle boundary artifacts. The first reason is likely stronger, as the localization accuracy depends on the size of the manipulated area, which is quantified below.
@@ 3.5 Implementation details
We used AUTOMATIC1111’s Stable Diffusion Web UI [1] for our experiments. Besides graphical interface with many plugins, it also provides a convenient batch processing.
To generate our dataset (RV5.1 and SD1.4), we used dynamic prompt [13]. This is the Web UI extension that implements an expressive template language for the generation of random or combinatorial prompts. In particular, we used the following prompt to get diversity in our datasat: “RAW photo, {older $|$ younger} {man $|$ woman $|$ lady $|$ girl $|$ boy} { {smiling $|$ staring} $|$ with glasses $|$ with hat $|$ with {brown $|$ blonde $|$ dark} {straight $|$ curly $|$ short} hair }, high quality portrait taken with Nikon camera, in {nature $|$ a city $|$ a room $|$ an office $|$ a park $|$ a street $|$ a forest }”.
All models were trained with PyTorch framework.